FEL Upgrade Project Weekly Brief
February 4-8, 2002


Late last week the optics team demonstrated that our proposed scheme for radius of curvature control of the primary optical cavity mirrors for the 10 kW Upgrade works extremely well in terms of resolution, linearity and concentricity. Details are given in the Optics Section below.

We continued to work up the cost tables and statements of work for the revisions to the ONR/DOE MOA that would encompass new FY02 and FY03 activities.

On Friday a visit was made to Northrop Grumman-Newport News to discuss potential collaboration on the proposed 100 kW Upgrade.

On Wednesday we had a visit from representatives from Gas Technology Institute who were interested in possible applications of the FEL for gas and oil extraction.

WBS 3 (Beam Physics):
Work to clarify diagnostic and correction system configurations continues. In support of this, and commissioning plans, the Demo machine model is being upgraded to describe, well what else - the Upgrade.

WBS 4 (Injector):
The support tube has been polished and a fixture is being manufactured to weld the internal mount ring prior to implantation. Fit up of the internal parts for the ball is being done prior to polishing the ball cathode. A mount for implanting the ball is being designed. A GaAs wafer was H2 cleaned for use in a field emission study. This is of interest for our quantification of cathode dark currents as we raise the gun voltage from 350kV in the 1 kW Demo to 500 kV in the Upgrade.

Gun HVPS & Tank - A cable tray, desk, table, and rack were moved out of the way in preparation for removing the Gun HVPS Tank by N. Wilson. The new tank is in the procurement process with a vendor not yet selected. Additional resistors were made this week to check the curing procedure and the voltage sensitivity of the material. Resistance measurements will be made again next week in an oil bath. Protective coatings will be applied next week also.

WBS 6 (RF):
Quarter HVPS - Except for the HV capacitors and the PLC expansion units, all major components have been installed in each HVPS. The schematics are at the 98% level and the wiring has started.

Quarter Klystrons - The design for the klystron water manifold is finished and drawings are being updated. Two additional klystron carts will be constructed as soon as the drawings are finished.

WBS 8 (Instrumentation):
Cable removal in the vault continues. We expect to have all of the cables we plan to remove out by the end of February and have begun plans to map magnet trim cables and power supply cables. Installation to begin around March 1.

The Machine Protection System input card design is complete and in to EECAD for final documentation prior to fabrication.

The technical requirements to shift the current 4 channel Beam Loss Monitoring system from a CAMAC to a VME based system are being determined and a plan is in progress to pursue this change.

New security features have been implemented to the FEL informational database. It can be accessed at: http://laser.jlab.org/data. Users with a JLAB e-mail address are permitted to log on immediately. Those users that do not have a JLAB e-mail address will be directed on the actions to take to become registered users of the system.

The drawings are being reviewed for the VME input card and fabrication is expected to begin soon. Drawings are also being revised for new beamviewer camera lamp rings to provide more light to the camera.

The machine shop has developed a device to put registration mark the foils on the beamviewers so when it is being used we will be better able to determine spot size and location on the viewer. New cameras are being built and the existing ones tested to support re-using them on the new machine.

Stepper motor control work continues. The chassis design is complete and fabrication is about to begin. The associated software is being developed.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o At Magnet Enterprises International in Oakland CA. the coil was potted, and cured and they are waiting for cool down. They have started winding additional coils.
o Core parts continue in rough machining, awaiting heat treatment of the entire batch.
Injector Dipoles (DU/DV)
o Proposals were received from a number of vendors and were evaluated. Questions were sent back whose answers should clarify their proposals.
o DULY Research is nearing completion of their checking the quality of the field of the Small Injector Dipole (GV).
Arc Dipoles (GY, GX, GQ)
o We started the electronic procurement of these magnets this week.
o Advanced Energy Systems (AES) is nearing completion of their analysis of the bend magnets at low excitation for low energy.
o Wang NMR of CA is continuing to insulate the conductor for the 180° Dipole coils and design of the winding form and potting mold.
o A subcontractor for Bosma Machine has torch cut out the GY core pieces to their circular shapes with the most beautiful burn surfaces I have ever seen. The process was done under water. In the mean time they are ordering hardware and are programming the machine tools for the automated machining operation.
3 inch quad (QX)
o Using our coordinate measuring machine, we found a discrepancy in the outside surface of the first 25 quadrupoles. It resulted in a re-shaving of the inner surfaces to remove a .001 error in the pole form at one coordinate caused by an error in our drawing. If we receive the remainder of the production run ion time, we are thinking that we will relegate the 25 early magnets to the UV line (26 quads required) to avoid magnet test stand, survey and alignment confusion.
o Additional measurements on a random magnet show that the pole surfaces are within .0005 inch of ideal? very good. Magnetic measurements on two quadrupoles indicate the fields are within specification. Dimensions to the fiducial surfaces are within specification but have to be accommodated at the magnet test stand and during survey and alignment. We are "working" those issues.
Trim Quad (QT)
o Milhous Control of Virginia has completed plating the core plates and is ready to start winding.
Sextupole (SF)
o DULY continues the final design at fine mesh.
Octupole (OT)
o Work on this magnet is on furlough until the sextupole is designed.
Corrector Dipoles (DB, DJ, GC BH)
o No news on qualification of the new sets of Standard Corrector Dipoles (DB/DJ). QX Quadrupole has the priority on the test stand.
o No news on the intended simple new vertical corrector for the Arcs (GC).
Beam Line and Vacuum
o AES continues drawing up the chambers through out the arcs. They have finished the finite element analysis of the most problematic chamber, the chamber at the GX in the optical cavity and have been told to stop further analysis because of the assurances that we are within the appropriate stress and deflection values from the initial analysis.
o We are finishing of the regions before the first arc and after the second arc.
o We took a final view at the layout of the extraction chicane where there was some interference in the stands upstream and some excess beam element length upstream. We decided to move the move the chicane upstream by about 5 cm.

WBS 10 (Wiggler):
Final commissioning data was taken on the second wiggler of the optical klystron this week. This data gives us the data we need for control scripts for the wiggler power supply. The data is being analyzed now. We hope to start commissioning data collection on the dispersion section next week.

WBS 11 (Optics):
The highlight of our week was a test of the 3" HR deformable mirror assembly (DMA) in the mirror test stand (MTS). Friday's experiments were done without using a laser to induce a thermal deformation on the surface. More experiments were done on the 5th, where we loaded the front surface with the output from a multimode Nd:YAG laser. To date, analyses to date reveal the following:

With or without surface heating, the radius of curvature (ROC) can be changed linearly by varying just one parameter. This will simplify our feedback control.

Our technique for inducing the change in the radius of curvature does not create any additional aberrations to the mirror surface. There is no indication of "print-through" of the cooling channels on the mirror surface.

With surface heating of the mirror at levels above that expected in the IR Upgrade, we can control the ROC over the range we require (several meters change from a starting ROC of 17.5m).

Still under investigation is to what degree the ROC control during surface heating mitigates the thermally-induced spherical aberration, particularly as a function of input heat flux. This is to further benchmark the predictions from our earlier work, as well as from the FEA modeling done last year by AES.

We are very pleased with the results to date, and wish to acknowledge the efforts of AES and our Mechanical Engineering group in designing and fabricating the DMA, and to the I&C group as well for their assistance in terminating the instrument and heater cabling.

Engineering checks of the revised drawing package for the optical cavity assemblies caught a few more changes that are being incorporated; this will delay release of the bid package by one week. The revised drawings for the ultraviewer mirror mount were signed Friday, and are being fabricated. (Thanks to K. Capek for disassembling the first article in order to retrieve parts). Questions on the OCMMS design were answered without any need to change the drawings; we will release them in a bit over a week. We held a teleconference with a potential vendor for the high power (50kW) laser beam dump, as well as for the backplane cooled insertable mirrors, we intend to send them some of our conceptual drawings. Design of the insertable mirror assemblies continues. We received a budgetary estimate on the leadscrews for the optical cavity assemblies, and plan to procure them next week.

We continue to procure parts to build a pulse compressor for the "waste" drive laser IR output.

Based on experiments done at one vendor's location, we have requested a budgetary estimate for the oscillator and timing electronics. This laser would initially be tested at 74.85 MHz, then reconfigured to operate at 748.5 MHz.